OpenStudy (littleheron):

WILL MEDAL AND FAN!! The point (5, -1) is a solution of the equation: y = 2x − 11 True False

OpenStudy (anonymous):

Do you have any clue on how to solve this problem?

OpenStudy (anonymous):

I literally only have 4 minutes before I need to go.

OpenStudy (littleheron):

nope

OpenStudy (anonymous):

Alright, coordinates are written as (x,y) So, you need to input these values into the equation and solve it. So, you have: -1=2(5)-11

OpenStudy (littleheron):

ok

OpenStudy (anonymous):

So, is it a solution?

OpenStudy (littleheron):

yes

OpenStudy (anonymous):

Yes, because 2*5=10 10-11=-1 -1=-1 So the answer is true.

OpenStudy (anonymous):

Do you understand how to solve the problem now?

OpenStudy (anonymous):

Alright, I got to get going, but good luck with the rest of your studying.

OpenStudy (littleheron):

Thanks